OpenStudy (anonymous):

How should I approach e^(2x) - e^x - 6 = 0 to solve for x?

myininaya (myininaya):

let u=e^x u^2-u-6=0 (u-3)(u+2)=0 u=3 => e^x=3 =>x=ln3 u=-2 cant have log of a negative x=ln3 is only solution

OpenStudy (anonymous):

thank you.